OpenStudy (anonymous):

Determine the Extreme values of the function f(x)=2 sin 4x + 3 XE[0, pi] I need help finding the x values for which their max/min using first derivitave

OpenStudy (anonymous):

screw the derivative do it in your head the largest sine can be is one the largest 2 times sine can be is 2 and the largest two times sine plus 3 can be is 5

OpenStudy (anonymous):

repeat for "smallest smallest sine can be is minus one smallest two times sine can be is minus two the smallest two times sine plus 3 can be is one

OpenStudy (anonymous):

yea, but im taking calculus, and my teacher expects to do the calculus way

OpenStudy (anonymous):

So, what is the derivative?

OpenStudy (anonymous):

tell your teacher that using derivatives in this case imparts no information the derivative of sine is cosine if you know about cosine, you know about sine you are chasing your tail

OpenStudy (anonymous):

As a side note, finding the critical points of the derivative is just as difficult as finding the minimum and maximum of the actual function here, since we know sin has max and min at 1 and 01.

OpenStudy (anonymous):

then the derivative of cosine is minus sine the derivative of minus sine is .... you get the idea it is stupid, a stupid waste of time going around in circles, saying you know about sine because you know about cosine and vice versa tell your math teacher (gently) that you are too smart for such nonsense and he/she should be as well

OpenStudy (anonymous):

the derivative of this function is 8 cos 4x,

OpenStudy (anonymous):

so it is set it equal to zero and solve for \(x\)

OpenStudy (anonymous):

once i do that i ge 45/2, and then i dont know what to do after that, i get confused

OpenStudy (anonymous):

\[8\cos(4x)=0\\ \cos(4x)=0\\ 4x=\frac{\pi}{2}\\ x=\frac{\pi}{8}\]

OpenStudy (anonymous):

you cannot use degrees in the context of calculus the trig functions are functions of numbers, not angles measured in degrees forget degrees

OpenStudy (anonymous):

btw you might as well have set \[2\sin(4x)=2\\ \sin(4x)=1\\ 4x=\frac{\pi}{2}\\ x=\frac{\pi}{8}\]same thing exactly

OpenStudy (anonymous):

ah i c makes sense now, thanks man

OpenStudy (anonymous):

Ummm, what about \(5\pi/8\)?

OpenStudy (anonymous):

I think \(3\pi/8\) is also a critical point.

OpenStudy (anonymous):

And also \(7\pi/8\).

OpenStudy (anonymous):

how did u get those?

OpenStudy (anonymous):

Basically, \(x\in [0,\pi] \implies 4x\in[0,4\pi]\)

OpenStudy (anonymous):

Then I got \(4x=\pi/2, \quad 4x=3\pi/2, \quad 4x=5\pi/2, \quad4x= 7\pi/2\).

OpenStudy (anonymous):

The fact that \(4\pi\) is 2 full rotations and there are 2 answers per rotation means there had to be more than one answer.

OpenStudy (anonymous):

Finally, \(0\) and \(2\pi\) are technically critical points simply for being on the boundaries.

OpenStudy (anonymous):

whoops, I mean \(0\) and \(\pi\).

OpenStudy (anonymous):

I do not get how you got \[3\pi/2\] and so on. im kind of confused on that part

OpenStudy (anonymous):

ooh wait ive been doing it wrong lol, i was doing 270/ 360 instead of doing 270/180 to get i into rads lol